OpenStudy (anonymous):

Consider the functions f(x) = x2 − 13 and g(x) = x + 5. What is the value of (f − g)(−4)?

OpenStudy (kainui):

f(x)-g(x) looks like (x2-13)-(x+5). Plug in -4 for x and you will find your answer.

OpenStudy (asnaseer):

@Kainui is correct. you are given:\[f(x)=x^2-13\]\[g(x)=x+5\]and are asked to compute "f-g" when x=-4. there are two ways of doing this. you can either first find f(-4) and g(-4) and then substract one from other like so:\[f(-4)=(-4)^2-13=16-13=3\]\[g(-4)=-4+5=1\]\[f(-4)-g(-4)=3-1=2\]or you can create a new function that represents "f-g", call it "h", and work out h(-4) as follows:\[h(x)=f(x)-g(x)=(x^2-13)-(x+5)=x^2-13-x-5=x^2-x-18\]\[h(-4)=(-4)^2-(-4)-18=16+4-18=2\]